# Perception Stitching: Zero-Shot Perception Encoder Transfer for Visuomotor Robot Policies

### v1

@@ 3.3 Implementation Details
Each input image is a $(84,84,3)$ RGB image. The visual encoder consists of a ResNet-18 network [25], followed by a spatial-softmax layer [17, 55], and then a 256-dimensional last layer at the module interface. We use 256 anchors to match the dimension of the latent representation. The eight-dimensional proprioception consists of the end effector position (3D), end effector quaternion (4D), and the gripper open width (1D). It is embedded by a 64-dimensional linear layer and then concatenated with two latent representations of the two images. We use the RNN-based action decoder for the Can and Door Open task and the MLP-based action decoder for other tasks. We list all the parameters of the neural network in the Appendix A with our code base. The actions are output by a Gaussian Mixture Model in the last layer of the policy network where the eight-dimensional action vector is the delta value of the $8$D proprioception. Before the perception stitching and zero-shot transfer, all policies are trained to 100% success rates with BC [65].
@@ 4 Experiments
Our experiments aim to (1) evaluate the effectiveness of PeS on enhancing zero-shot policy transfer on the stitched policies and (2) understand the mechanisms behind the performance advantage of PeS. Our experiments consist of four parts. First, we evaluate the zero-shot transfer performance of PeS in five simulated manipulation tasks with six variations of camera configurations. We then assess the performance of PeS in four real-world manipulation tasks with three different camera configurations. In addition, we analyze the latent representations of the visual encoders both quantitatively and qualitatively to understand the effectiveness of PeS in latent space alignment. Lastly, we adopt Gradient-weighted Class Activation Mapping (Grad-CAM) [74] to visualize the regions that the visuomotor policies focus on, offering an intuitive insight into the success of PeS.
@@
Figure 4: Simulation Experiment Setup. (a) Five simulation tasks from Robomimic [55] benchmark. (b) Six camera configuration variations include. (c) Four camera mounting positions.
4.1 Simulation Experiment Setup and Baselines
@@
Evaluation Tasks We evaluate PeS in five manipulation tasks from the Robomimic [55] benchmark (Fig. 4(a)): (1) Push: a robot starts from a random initial position in the air and pushes the cube forward for a certain distance (30 cm in simulation, 20 cm in real world). (2) Lift: a robot lifts up a cube randomly placed on the table. (3) Can task: a robot picks up a can on its right and places it into the bins on its left. (4) Stack: a robot picks up the red cube on the table and stacks it on the green cube. (5) Door Open: a robot grabs the handle and pulls the door open.
@@
For each task, we test PeS with six camera configurations in Fig. 4(b). (1) Masked: A black square mask is put on the upper-left corner of the image to mimic a partially occluded camera lens. Although the mask makes little difference for humans, our experiments suggest a strong negative effect on current visuomotor policies. (2) Zoom in: zoom in the camera to enlarge the object in a smaller field of view. 3) Blurred: Gaussian blur effect is added to the image. (4) Gaussian Noise: Gaussian noise is added to the image to simulate low-light condition image capture. (5) Fisheye: a fisheye effect is simulated. (6) Camera Position: the camera is put in one out four possible positions as in Fig. 4(c).
@@
Camera Configurations The first five configurations used in-hand and close-up cameras. For example, in the Push-Blurred experiment, we first train policy $1$ with a normal in-hand camera and a blurry close-up camera, and policy $2$ with a blurry in-hand camera and a normal close-up camera. Then we stitch the close-up view visual encoder of the policy $2$ to the in-hand view visual encoder and action decoder of the policy $1$ and test this stitched policy in the Push task with a normal in-hand camera and a normal close-up camera without any fine-tuning. The last configuration is designed to test zero-shot transfer ability across different camera positions. Policy $1$ is trained with an in-hand camera and a front-view camera. Policy $2$ is trained with a close-up view camera and a side-view camera. Then the side-view encoder of policy $2$ is stitched to the in-hand view encoder and action decoder of policy $1$, and tested with the in-hand and side-view cameras.
@@
Baselines We adopt three baseline methods and two ablation studies: (1) Devin et al. [12] uses a task network module to embed the observation (task state) agnostic to the robot itself, concatenates it with the proprioceptive robot state, and then inputs it into the downstream robot network module. For a fair comparison, the task module and the robot module use the same structure as our PeS method. The key difference is that Devin et al. [12] does not adopt the relative representation and does not apply our disentanglement regularization. (2) Cannistraci et al. [6] (linear) uses four similarity functions, $L_{1}$, $L_{2}$, $L_{\infty}$, $cos$ to calculate four relative representations and then calculates the element-wise linear sum of the four relative representations to be the final latent state passed over to the action decoder. The weights of the $L_{1}$, $L_{2}$, $L_{\infty}$, $cos$ similarities are $0.05$, $0.1$, $0.05$, and $0.8$. In contrast, our PeS method only uses the cosine similarity. (3) Cannistraci et al. [6] (nonlinear) also adopts these four similarity functions but passes them through a non-linear layer ( LayerNorm, a Linear layer, and a Tanh) followed by an element-wise summation. (4) The PeS (w/o disent. loss) ablation method does not have the disentanglement loss Eq. 3. (5) The PeS (w. l1 & l2 loss) replaces the disentanglement loss in PeS with L1 and L2 penalty, as shown in Eq. 3.
@@
Results Tab. 1 shows that PeS and its two ablation methods can all achieve satisfying performance with around 90% of success rates on Push and Lift. Cannistraci et al. 2024 (linear) has lower success rates around 70% - 80%, suggesting that introducing a linear combination of multiple similarity measurements for the relative representation hurts the model’s performance. Since isometric transformation is the major transformation observed at the latent space, the single cosine similarity of PeS offers simple but effective features. Cannistraci et al. 2024 (non-linear) performs worse than the linear baseline and reaches about 20% to 50% success rates. We believe that the nonlinear combination introduces learnable parameters for the non-linear summation process. However, these parameters are not optimized to align the latent spaces but are optimized for the accuracy of behavior cloning, leading to unaligned latent space and hurting the performance. Devin et al. 2017 baseline achieves very low success rates between 18% to 33%, indicating that the unaligned latent representations cause the failure of the stitched policy for zero-shot transfer.

### v2

@@ 3.3 Implementation Details
Each input image is a $(84,84,3)$ RGB image. The visual encoder consists of a ResNet-18 network (He et al., 2016), followed by a spatial-softmax layer (Finn et al., 2016; Mandlekar et al., 2021), and then a 256-dimensional last layer at the module interface. We use 256 anchors to match the dimension of the latent representation. The eight-dimensional proprioception consists of the end effector position (3D), end effector quaternion (4D), and the gripper open width (1D). It is embedded by a 64-dimensional linear layer and then concatenated with two latent representations of the two images. We use the RNN-based action decoder for the Can and Door Open task and the MLP-based action decoder for other tasks. We list all the parameters of the neural network in the Appendix A with our code base. The actions are output by a Gaussian Mixture Model in the last layer of the policy network where the eight-dimensional action vector is the delta value of the $8$D proprioception. Before the perception stitching and zero-shot transfer, all policies are trained to 100% success rates with BC (Pomerleau, 1988). The dataset of each task contains $200$ trajectories of the expert demonstrations. The pseudo code of PeS is presented in Appendix B B.
@@ 4 Experiments
Our experiments aim to (1) evaluate the effectiveness of PeS on enhancing zero-shot policy transfer on the stitched policies and (2) understand the mechanisms behind the performance advantage of PeS. Our experiments consist of five parts. First, we evaluate the zero-shot transfer performance of PeS for training double-camera-view visuomotor policies in five simulated manipulation tasks with seven variations of camera configurations. Next, we apply PeS to single-camera-view and triple-camera-view policies to test its generalizability to arbitrary camera number settings. We then assess the performance of PeS in four real-world manipulation tasks with three different camera configurations. In addition, we analyze the latent representations of the visual encoders both quantitatively and qualitatively to understand the effectiveness of PeS in latent space alignment. Lastly, we adopt Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2017) to visualize the regions that the visuomotor policies focus on, offering an intuitive insight into the success of PeS.
@@
Figure 4: Simulation Experiment Setup. (a) Five simulation tasks from Robomimic (Mandlekar et al., 2021) benchmark. (b) Seven camera configuration variations include. (c) Four camera mounting positions.
4.1 Simulation Experiments with Double Camera Views
@@
Evaluation Tasks We evaluate PeS in five manipulation tasks from the Robomimic (Mandlekar et al., 2021) benchmark (Fig. 4(a)): (1) Push: a robot starts from a random initial position in the air and pushes the cube forward for a certain distance (30 cm in simulation, 20 cm in real world). (2) Lift: a robot lifts up a cube randomly placed on the table. (3) Can task: a robot picks up a can on its right and places it into the bins on its left. (4) Stack: a robot picks up the red cube on the table and stacks it on the green cube. (5) Door Open: a robot grabs the handle and pulls the door open.
@@
Camera Configurations For each task, we test PeS with seven camera configurations in Fig. 4(b). (1) Masked: A black square mask is put on the upper-left corner of the image to mimic a partially occluded camera lens. Although the mask makes little difference for humans, our experiments suggest a strong negative effect on current visuomotor policies. (2) Zoom in: zoom in the camera to enlarge the object in a smaller field of view. 3) Blurred: Gaussian blur effect is added to the image. (4) Gaussian Noise: Gaussian noise is added to the image to simulate low-light condition image capture. (5) Fisheye: a fisheye effect is simulated. (6) Lighting: the image is darker or brighter than the normal images, mimicking the overexposed or underexposed situation. (7) Position: the camera is put in one out of four possible positions as in Fig. 4(c).
@@
The first six configurations used in-hand and close-up cameras. For example, in the Push-Blurred experiment, we first train policy $1$ with a normal in-hand camera and a blurry close-up camera, and policy $2$ with a blurry in-hand camera and a normal close-up camera. Then we stitch the close-up view visual encoder of the policy $2$ to the in-hand view visual encoder and action decoder of the policy $1$ and test this stitched policy in the Push task with a normal in-hand camera and a normal close-up camera without any fine-tuning. The last configuration is designed to test zero-shot transfer ability across different camera positions. Policy $1$ is trained with an in-hand camera and a front-view camera. Policy $2$ is trained with a close-up view camera and a side-view camera. Then the side-view encoder of policy $2$ is stitched to the in-hand view encoder and action decoder of policy $1$, and tested with the in-hand and side-view cameras.
@@
Baselines We adopt four baseline methods from previous works and two ablation studies: (1) RT-1 (Brohan et al., 2022) is a multi-task large model for robotic control built on the Transformer architecture. It is trained on a large dataset of various robotic tasks and can be generalized in zero-shot to new tasks. (2) Devin et al. (2017) uses a task network module to embed the observation (task state) agnostic to the robot itself, concatenates it with the proprioceptive robot state, and then inputs it into the downstream robot network module. For a fair comparison, the task module and the robot module use the same structure as our PeS method. The key difference is that Devin et al. (2017) does not adopt the relative representation and does not apply our disentanglement regularization. (3) Cannistraci et al. (2023) (linear) uses four similarity functions, $L_{1}$, $L_{2}$, $L_{\infty}$, $cos$ to calculate four relative representations and then calculates the element-wise linear sum of the four relative representations to be the final latent state passed over to the action decoder. The weights of the $L_{1}$, $L_{2}$, $L_{\infty}$, $cos$ similarities are $0.05$, $0.1$, $0.05$, and $0.8$. In contrast, our PeS method only uses cosine similarity. (4) Cannistraci et al. (2023) (nonlinear) also adopts these four similarity functions but passes them through a non-linear layer (consists of a LayerNorm, a linear layer, and a Tanh) followed by an element-wise summation. (5) The PeS (w/o disent. loss) ablation method does not have the disentanglement loss Eq. 7. (6) The PeS (w. l1 & l2 loss) replaces the disentanglement loss in PeS with L1 and L2 penalty, as shown in Eq. 8. In addition, we also report the performance of policies 1 and 2 to be stitched if they are transferred to the target environment in zero-shot. These two baselines help us understand how much this transfer preserves the performance of the policy in the initial setting, and how much extra performance increase PeS can bring.
Mask Zoom in Blurred Noise Fisheye Position Lighting Average
Push Policy 1 94.0$\pm$2.83 88.7$\pm$0.94 88.0$\pm$1.63 74.7$\pm$5.73 69.3$\pm$0.94 41.3$\pm$0.94 38.0$\pm$1.63 70.6
Policy 2 96.7$\pm$0.94 86.0$\pm$1.63 84.7$\pm$3.40 73.3$\pm$2.49 68.7$\pm$1.88 7.3$\pm$0.97 36.0$\pm$2.83 64.7
RT-1 95.3$\pm$2.49 89.3$\pm$7.36 86.0$\pm$4.32 77.3$\pm$8.22 76.0$\pm$8.49 45.3$\pm$5.73 40.0$\pm$6.53 72.7
Devin et al. 2017 60.7$\pm$10.6 8.7$\pm$4.99 16.7$\pm$3.77 59.3$\pm$6.80 29.3$\pm$7.36 19.3$\pm$5.73 36.7$\pm$3.40 33.0
Cannistraci et al. 2024 (linear) 89.3$\pm$4.11 94.0$\pm$2.83 64.7$\pm$1.89 74.7$\pm$6.18 74.0$\pm$2.83 78.7$\pm$2.49 87.3$\pm$0.94 80.4
Cannistraci et al. 2024 (non-linear) 12.7$\pm$1.89 18.7$\pm$4.99 42.8$\pm$3.27 23.3$\pm$0.94 6.0$\pm$4.32 5.3$\pm$2.49 32.7$\pm$3.40 20.2
PeS (w/o disent. loss) 100.0$\pm$0.0 86.0$\pm$2.83 80.7$\pm$9.84 100.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 86.7$\pm$0.94 93.3
PeS (w. l1 & l2 loss) 88.7$\pm$4.99 95.3$\pm$1.89 90.0$\pm$5.66 100.0$\pm$0.0 93.3$\pm$0.94 80.7$\pm$4.99 89.3$\pm$0.94 91.0
PeS 100.0$\pm$0.00 100.0$\pm$0.00 95.3$\pm$0.94 100.0$\pm$0.0 92.7$\pm$2.50 100.0$\pm$0.00 94.0$\pm$4.32 97.4
Lift Policy 1 86.0$\pm$1.63 24.7$\pm$3.77 68.0$\pm$4.90 88.0$\pm$1.63 38.7$\pm$5.25 0.0$\pm$0.00 15.3$\pm$6.18 45.8
Policy 2 82.0$\pm$4.32 12.0$\pm$7.12 91.3$\pm$3.40 83.3$\pm$0.94 13.3$\pm$0.94 0.0$\pm$0.00 44.0$\pm$2.83 46.6
RT-1 86.0$\pm$2.82 0.0$\pm$0.00 86.7$\pm$6.24 85.3$\pm$4.99 14.0$\pm$3.27 35.0$\pm$7.07 46.0$\pm$4.32 50.4
Devin et al. 2017 0.0$\pm$0.00 5.3$\pm$2.49 48.0$\pm$5.89 9.3$\pm$4.11 14.7$\pm$4.99 36.0$\pm$1.63 18.7$\pm$4.99 18.9
Cannistraci et al. 2024 (linear) 72.7$\pm$3.77 64.0$\pm$2.83 86.0$\pm$4.32 68.7$\pm$1.88 88.7$\pm$1.88 57.3$\pm$2.49 56.0$\pm$5.89 70.5
Cannistraci et al. 2024 (non-linear) 89.3$\pm$2.49 36.0$\pm$3.27 52.7$\pm$3.40 93.3$\pm$2.49 16.7$\pm$2.49 21.3$\pm$0.94 37.3$\pm$2.49 49.5
PeS (w/o disent. loss) 83.3$\pm$6.60 80.7$\pm$5.73 93.3$\pm$0.94 91.3$\pm$5.73 79.3$\pm$2.49 93.3$\pm$2.49 76.7$\pm$3.40 85.4
PeS (w. l1 & l2 loss) 97.3$\pm$2.49 85.3$\pm$0.94 90.7$\pm$0.94 86.0$\pm$4.32 88.0$\pm$1.63 84.7$\pm$3.77 64.7$\pm$3.40 85.2
PeS 92.7$\pm$2.50 94.7$\pm$1.89 89.3$\pm$4.11 96.0$\pm$1.63 88.7$\pm$0.94 93.0$\pm$0.03 84.7$\pm$6.60 91.3
Table 1: Zero-Shot Transfer Success Rates in basic Simulation tasks. We test all methods across six different visual configurations and also calculate the average performance. In the two basic tasks, Push and Lift, PeS and its two ablation methods can all get about a 90% success rate on average. The Cannistraci et al. 2024 (linear) baseline has 70% - 80% of success rate, while the Cannistraci et al. 2024 (non-linear) and Devin et al. 2017 baselines perform poorly with about 20% - 50% of success rate.
@@
Results Tab. 1 shows that PeS and its two ablation methods can all achieve satisfying performance with around 90% of success rates on Push and Lift. Compared with directly transferring policy 1 or policy 2 to the new environment, PeS results in a 25% to 45% success rate increase. RT-1 demonstrates satisfying robustness on moderate visual variations such as “Mask” and “Noise”. However, RT-1 doesn’t perform well when the task becomes harder and the visual configuration changes more drastically. For example, it only achieves a 14% of success rate in the Lift-Fisheye task. Cannistraci et al. 2024 (linear) has lower success rates around 70% - 80%, suggesting that introducing a linear combination of multiple similarity measurements for the relative representation hurts the model’s performance. Since isometric transformation is the major transformation observed at the latent space, the single cosine similarity of PeS offers simple but effective features. Cannistraci et al. 2024 (non-linear) performs worse than the linear baseline and reaches about 20% to 50% success rates. We hypothesize that the nonlinear combination introduces learnable parameters for the non-linear summation process. However, these parameters are not optimized to align the latent spaces but are optimized for the accuracy of behavior cloning, leading to unaligned latent space and hurting the performance. Devin et al. 2017 baseline achieves very low success rates between 18% to 33%, indicating that the unaligned latent representations cause the failure of the stitched policy for zero-shot transfer.
Mask Zoom in Blurred Noise Fisheye Position Lighting Average
Can Policy 1 76.7$\pm$3.40 12.7$\pm$0.94 80.0$\pm$2.83 83.3$\pm$2.49 16.0$\pm$1.63 0.0$\pm$0.00 0.0$\pm$0.00 38.4
Policy 2 80.0$\pm$0.00 4.0$\pm$1.63 78.0$\pm$0.00 86.7$\pm$0.94 13.3$\pm$4.11 0.0$\pm$0.00 2.0$\pm$0.00 37.7
RT-1 88.3$\pm$6.24 0.0$\pm$0.00 86.7$\pm$8.50 88.3$\pm$2.36 0.0$\pm$0.00 8.3$\pm$6.24 4.0$\pm$0.00 39.4
Devin et al. 2017 19.3$\pm$5.25 24.7$\pm$1.89 2.7$\pm$1.89 6.0$\pm$4.32 29.3$\pm$3.40 1.3$\pm$1.89 3.3$\pm$1.89 12.4
Cannistraci et al. 2024 (linear) 33.3$\pm$0.94 48.0$\pm$1.63 48.7$\pm$2.49 65.3$\pm$0.94 26.7$\pm$3.77 34.7$\pm$3.77 42.7$\pm$0.94 42.8
Cannistraci et al. 2024 (non-linear) 72.7$\pm$0.94 24.7$\pm$2.49 37.3$\pm$4.99 42.7$\pm$3.40 8.7$\pm$1.89 39.3$\pm$1.89 38.7$\pm$1.89 37.7
PeS (w/o disent. loss) 44.7$\pm$8.06 89.3$\pm$4.11 34.7$\pm$4.11 30.7$\pm$6.80 92.7$\pm$2.50 44.7$\pm$3.40 42.7$\pm$0.94 54.2
PeS (w. l1 & l2 loss) 47.3$\pm$0.94 58.7$\pm$1.88 54.0$\pm$8.64 36.0$\pm$7.12 58.7$\pm$1.88 64.7$\pm$6.60 38.0$\pm$4.32 51.1
PeS 83.3$\pm$5.24 89.3$\pm$2.49 74.0$\pm$2.83 78.7$\pm$4.11 56.0$\pm$2.83 78.7$\pm$2.49 73.3$\pm$6.80 76.2
Stack Policy 1 66.0$\pm$2.83 16.0$\pm$5.89 1.3$\pm$0.94 65.3$\pm$0.94 26.7$\pm$2.49 0.0$\pm$0.00 15.3$\pm$6.18 27.2
Policy 2 83.3$\pm$2.49 0.0$\pm$0.00 4.0$\pm$1.63 64.0$\pm$0.00 2.0$\pm$1.63 2.0$\pm$1.63 24.0$\pm$2.83 25.6
RT-1 85.0$\pm$4.08 0.0$\pm$0.00 0.0$\pm$0.00 76.7$\pm$4.71 13.3$\pm$5.73 0.0$\pm$0.00 31.3$\pm$0.94 29.5
Devin et al. 2017 0.7$\pm$0.94 8.0$\pm$1.63 0.7$\pm$0.94 24.0$\pm$2.83 0.0$\pm$0.00 14.0$\pm$3.27 4.7$\pm$2.49 7.4
Cannistraci et al. 2024 (linear) 47.3$\pm$0.94 62.0$\pm$4.32 32.7$\pm$3.77 30.7$\pm$0.94 54.0$\pm$8.64 14.7$\pm$6.18 0.7$\pm$0.94 34.6
Cannistraci et al. 2024 (non-linear) 10.0$\pm$1.63 12.0$\pm$0.00 0.0$\pm$0.00 3.3$\pm$0.94 0.0$\pm$0.00 0.7$\pm$0.94 8.7$\pm$3.77 5.0
PeS (w/o disent. loss) 34.0$\pm$11.43 10.7$\pm$4.11 62.0$\pm$10.71 34.0$\pm$7.12 22.7$\pm$3.77 26.0$\pm$4.32 36.7$\pm$8.06 32.3
PeS (w. l1 & l2 loss) 92.7$\pm$0.94 98.0$\pm$0.00 62.7$\pm$6.60 24.0$\pm$4.90 59.3$\pm$7.36 58.7$\pm$1.88 48.0$\pm$4.32 63.3
PeS 94.7$\pm$0.94 96.7$\pm$0.94 90.0$\pm$1.63 96.7$\pm$1.89 97.3$\pm$2.49 80.0$\pm$4.90 82.0$\pm$1.63 91.1
Door Open Policy 1 95.3$\pm$1.89 0.0$\pm$0.00 19.3$\pm$4.71 66.0$\pm$4.32 4.0$\pm$1.63 0.0$\pm$0.00 0.7$\pm$0.94 26.5
Policy 2 89.3$\pm$0.94 5.3$\pm$1.89 21.3$\pm$5.25 52.7$\pm$0.94 4.0$\pm$1.63 0.0$\pm$0.00 14.7$\pm$3.40 26.8
RT-1 96.0$\pm$1.63 0.0$\pm$0.00 26.7$\pm$3.40 94.7$\pm$3.40 0.7$\pm$0.94 0.0$\pm$0.00 0.0$\pm$0.00 31.2
Devin et al. 2017 9.3$\pm$4.11 5.3$\pm$0.94 0.0$\pm$0.00 4.0$\pm$1.63 0.7$\pm$0.94 0.0$\pm$0.00 6.0$\pm$1.63 3.6
Cannistraci et al. 2024 (linear) 0.0$\pm$0.00 1.3$\pm$0.94 10.7$\pm$2.49 10.7$\pm$4.99 2.0$\pm$1.63 47.3$\pm$9.29 12.0$\pm$7.12 12.0
Cannistraci et al. 2024 (non-linear) 26.0$\pm$2.83 31.3$\pm$4.99 49.3$\pm$8.22 48.0$\pm$5.89 62.7$\pm$3.40 44.7$\pm$3.40 33.3$\pm$4.99 42.2
PeS (w/o disent. loss) 24.7$\pm$7.71 44.0$\pm$2.83 34.7$\pm$3.77 0.7$\pm$0.94 36.7$\pm$0.94 23.3$\pm$3.40 24.0$\pm$2.83 26.9
PeS (w. l1 & l2 loss) 4.0$\pm$1.63 78.0$\pm$5.66 3.3$\pm$0.94 2.0$\pm$1.63 42.7$\pm$4.99 6.0$\pm$3.26 14.7$\pm$4.99 21.5
PeS 58.7$\pm$4.11 68.7$\pm$0.94 70.7$\pm$0.94 52.7$\pm$3.40 64.7$\pm$4.99 48.7$\pm$3.40 56.7$\pm$5.73 60.1
Table 2: Zero-Shot Transfer Success Rates in difficult Simulation tasks. The three difficult tasks include Can, Stack, and Door Open. We test all methods across six different visual configurations and also calculate the average performance. The three baselines achieve about 0% - 40% of success rates in these tasks. In the Can and Stack tasks, PeS achieves 75% - 95% of success rates, while the ablation methods achieves about 30% - 65% of success rates. In the Door task, PeS achieves 60.7% of success rates, while the two ablation methods only have 20% - 30% of success rates.
Tab. 2 reports the performance in more difficult tasks: Can, Stack, and Door Open. PeS achieves 60% to 93% of success rates, which has about 30% to 60% of advantage to the base policy 1 and 2. The success rates of all four baselines are lower than 43%. This result suggests that previous methods cannot solve the zero-shot transfer problem for difficult tasks with satisfactory performance, which usually involve complicated visual background (e.g., Can), long-horizon multi-stage motions (e.g., Stack), and articulated objects (e.g., Door Open). Among these baselines, we find that RT-1 sometimes can achieve very good performance when the visual variation is moderate. For example, in the Can-Mask task and the Door Open-Noise task, RT-1 achieves better performance than all other methods including PeS. This is somewhat not surprising since RT-1 was trained on a much larger dataset with a much larger model that could cover some of these variations in training. However, for the harder tasks with more drastic visual configuration changes, such as “Door-Fisheye” and “Stack-Camera Position”, we found that the RT-1 method could not perform well. This additional result suggests that the trade-off between modular and end-to-end architectures is that an end-to-end network is simpler in the end-to-end training procedure but usually cannot generalize to drastic changes in visual configuration that have not been encountered during training. In contrast, although the modular policies usually involve more careful structure designs in the training procedure, our results show that they performed better in drastic visual changes.
We also notice that the two ablation methods have significant performance drops (20% to 60%) in these difficult tasks, showing that the disentanglement regularization in PeS largely improves the zero-shot transfer performance. Although using L1 and L2 regularization also leads to good success rates in some cases, the disentanglement regularization has more consistent performance and higher success rates on average.
4.2 Simulation Experiments with Single Camera View and Triple Camera Views
While most visuomotor policies adopt two cameras in their applications, we also conduct experiments for policies with only one camera and policies with three cameras to verify the generalizability of PeS to arbitrary camera number settings.
Single Camera View We train two policies with two different cameras separately. Each policy still has two visual encoders, but one visual encoder takes in a black image and the other encoder takes in the images from the single camera. For both encoders of the policy, they use the same anchor images set collected by the only camera, because we assume that each policy is trained separately and only has access to the dataset of one camera.
In the perception stitching process, we stitch the visual encoder of policy $2$ to the visual encoder and action decoder of policy $1$, as shown in Figure 5, and assemble a stitched policy for double camera inputs. We suggest that this is the most reasonable setting for the perception stitching of single-camera policies. In contrast, if we remove the black image encoder and stitch the only vision encoder of the policy $1$ to the action decoder of the policy $2$, applying this reassembled policy in either environment $1$ or $2$ won’t lead to better performance compared with the original policy trained to a maximum success rate in that environment.
Ideally, the decoder should learn to completely ignore the latent representations from the encoder with empty input if we train the policy with a large amount of expert data and many epochs. In practice, however, the policy is trained with only $200$ expert demonstrations of data and limited epochs. The policy will learn to assign more importance to the encoder with RGB image input and less importance to the encoder with empty input, but it will not fully ignore that encoder with empty input. Therefore, stitching another encoder to replace the encoder with empty input will have less influence on the policy performance compared with the double camera experiments, but the disturbance caused by this new encoder will still exist due to the latent space misalignment, although in some easier experiments, the disturbance is marginal. As supported by our single-camera experiment results, the performance decreases of the baseline methods after the zero-shot transfer are smaller, compared with that in the double-camera experiments, but the performance decreases are not zero.
Push Lift Can Stack Door Open Average